OpenStudy (anonymous):

A plane landing at PDX has to descend 1650ft over a diagonal distance of 9700ft. What is the angle of depression of the plane? help would be nicce<33

OpenStudy (anonymous):

Let D for angle of depression sinD=opp/hy sinD=1650/9700 use caculator: \[ \sin^{-1}(\frac{1650}{9700})=9.79^{o}\]\[or: 10^{o}\]